OpenStudy (anonymous):

Samantha is going on vacation for the summer and is trying to choose between two different plans. The first plan costs $450 for 3 days at a hotel and 2 days at an amusement park. The second plan offers 5 days at the same hotel and 3 days at the amusement park for $700. The cost of 1 day at the hotel and the amusement park is the same under both plans. How much does a 1 day trip to the amusement park cost

OpenStudy (anonymous):

Let h = cost of hotel per night and a = cost of amusement park for one day. Then we have two equations as follows: $450 = 3h + 2a $700 = 5h + 3a multiply first eqn by 5 and second by -3 $2250 = 15h + 10 a -$2100 = -15h - 9a Now add them: $150 = a The cost of one day at the amusement park is $150. Hope that helps!

OpenStudy (anonymous):

thanks I understand!

OpenStudy (anonymous):

:)